# Towards Explainable Bilingual Multimodal Misinformation Detection and Localization

### v1

@@ Abstract
The increasing realism of multimodal content has made misinformation more subtle and more complex to detect, especially in news media, where images are frequently paired with bilingual subtitles (e.g., Chinese-English). Such content often includes localized image edits and cross-lingual inconsistencies that jointly distort meaning while remaining superficially plausible. We introduce BiMi, a bilingual multimodal framework that jointly performs region-level localization, cross-modal and cross-lingual consistency detection, and natural language explanation for misinformation analysis. To support generalization, BiMi integrates an online retrieval module that supplements model reasoning with up-to-date external context. We further release BiMiBench, a large-scale and comprehensive benchmark constructed by systematically editing real news images and subtitles, comprising 104,000 samples with realistic manipulations across visual and linguistic modalities. To enhance interpretability, we apply Group Relative Policy Optimization (GRPO) to improve explanation quality, marking the first use of GRPO in this domain. Extensive experiments demonstrate that BiMi outperforms strong baselines by up to +8.9 in classification accuracy, +15.9 in localization accuracy, and +2.5 in explanation BERTScore, advancing state-of-the-art performance in realistic, multilingual misinformation detection. Code, models, and datasets are avaliable at https://github.com/RaffeLegend/misinformation.
@@ 1 Introduction
The rapid advancement of generative models [1, 2, 3, 4, 5] has drastically lowered the barrier to producing highly realistic multimodal content [6]. While enabling creative applications, this progress has also escalated the proliferation of misinformation that intricately blends visual and textual elements [7, 8]. A particularly insidious form emerges in news media, where images are frequently overlaid with bilingual subtitles (e.g., Chinese-English). Here, subtle visual tampering and cross-lingual inconsistencies are exploited to distort facts while maintaining surface-level plausibility. This becomes especially dangerous for monolingual audiences, who may only comprehend one language and thus remain oblivious to manipulations in the other. A stark example occurred during early 2020, when a CDC report confirming the first U.S. case of COVID-19 community transmission [^1] https://edition.cnn.com/2020/02/27/health/us-cases-coronavirus-community-transmission/index.html was mistranslated by some Chinese social media as asserting that the virus originated in the United States. This misleading bilingual manipulation fueled significant public misunderstanding and geopolitical tensions. Such cases illustrate how multimodal misinformation increasingly leverages localized image edits and asymmetric subtitle translations to manipulate public perception across language communities. Addressing such sophisticated attacks demands more than conventional binary misinformation detection. It requires systems capable of region-level localization of manipulations, identification of cross-modal and cross-lingual inconsistencies, and generation of coherent, cross-lingual natural language explanations—capabilities that remain underexplored.
Despite encouraging progress [7, 8, 9, 10, 11], existing systems exhibit three critical limitations. First, most rely on global image-text matching, lacking the fine-grained region-level localization necessary to expose subtle manipulations. Second, their explanations are often superficial and generic, providing little actionable insight into why content is misleading. Third, they neglect the modeling of bilingual subtitle inconsistencies, overlooking the semantic divergence arising from intentionally misleading translations. These shortcomings underline the need for a unified framework that reasons at a fine granularity across visual and bilingual textual modalities.
@@
Figure 1: Comparison of traditional vs. bilingual misinformation detection tasks. Traditional tasks focus on visual-text consistency with limited outputs (left). Our setting uses tampered news images and bilingual subtitles, enabling richer outputs including region localization, cross-modal consistency, and explanation (right). Red indicates error, green indicates correctness. Best viewed in color.
@@
However, building such a framework is technically challenging. Region localization demands precise visual-textual alignment that is beyond the capacity of most pretrained MLLMs. Generating faithful explanations in bilingual, multimodal settings requires accurate identification and articulation of inconsistencies across modalities and languages. Moreover, cross-lingual reasoning introduces unique complexities, as even fluent translations may introduce subtle but impactful semantic shifts that obscure the misinformation signal. These challenges are further exacerbated in the news domain, characterized by rapidly evolving, diverse, and context-dependent content.
@@
To tackle these challenges, we introduce BiMi, the first bilingual multimodal misinformation detection framework that seamlessly integrates region-level manipulation localization, cross-modal and cross-lingual consistency detection, and natural language explanation generation. An overview of our task and outputs is illustrated in Figure 1. To enhance generalization to emerging events, BiMi is equipped with an online retrieval module that augments model reasoning with real-time external knowledge. Furthermore, to facilitate systematic evaluation, we construct BiMiBench, the first large-scale benchmark for this setting, comprising 104,000 samples with realistic manipulations of news images and bilingual subtitles. Through extensive experiments, BiMi achieves state-of-the-art performance, surpassing strong baselines by +8.9 in classification accuracy, +15.9 in localization accuracy, and +2.5 in explanation BERTScore.
Our contributions are summarized as follows:
@@
  • We release BiMiBench, the first large-scale benchmark for bilingual multimodal misinformation detection, containing 104K samples with fine-grained image manipulations and bilingual subtitle inconsistencies.
@@
  • We propose BiMi, a unified framework that detects image-subtitle misinformation via region localization, cross-modal consistency detection, and explanation generation, addressing the limitations of prior works.
@@
  • We introduce the first application of Group Relative Policy Optimization (GRPO) [12] in this domain, enhancing explanation quality in multimodal, bilingual scenarios.
@@
  • Our framework achieves new state-of-the-art performance on BiMiBench, substantially improving classification, localization, and explanation metrics over competitive baselines.
@@ 2.1 Misinformation Detection
Datasets. Early misinformation detection datasets can be grouped into three main categories. Text-only datasets: FEVER [13] and FakeNewsNet [14] focus exclusively on textual fact verification and lack any visual modality, limiting their applicability to multimodal scenarios. Basic multimodal datasets: Fakeddit [15], MAIM [16], and EMU [17] introduce image-text pairs, but are typically limited to monolingual (English) content, coarse visual edits, and offer no explicit supervision for localization or explanation. Semantically aligned multimodal datasets: More recent benchmarks such as NewsCLIPpings [18], COSMOS [19], MMFakeBench [8] and DGM4 [9] move toward finer-grained multimodal reasoning. However, they remain focused on caption-style inputs and overlook subtitle-level or cross-lingual inconsistencies.
To address these limitations, we introduce BiMiBench, a large-scale benchmark curated from real-world news content that captures both visual and textual manipulations across five fine-grained misinformation types. As shown in Table 2.1, BiMiBench is the only benchmark that simultaneously supports bilingual subtitles, region-level localization, and both visual and textual tampering, offering the richest annotation scope among existing datasets. To evaluate the quality of visual manipulations, Figure 2 compares BiMiBench with recent benchmarks using a perceptual realism analysis across seven dimensions (e.g., content, composition, color), demonstrating that BiMiBench achieves significantly higher realism and diversity.
Table 1: Comparison of misinformation datasets. BiMiBench uniquely supports manipulations based on bilingual subtitles and localized content.
Dataset Multi- Sub- Bi- Textual Visual Local- News Mask
modal title lingual Manip. Manip. ization Domain
FEVER [13]
FakeNewsNet [14]
Fakeddit [15]
MAIM [16]
EMU [17]
NewsCLIPpings [18]
COSMOS [19]
DGM4 [9]
MMFakeBench [8]
@@
Figure 2: Quality comparison of manipulated images across datasets, by the method of [20].
Methods. Current methods for multimodal misinformation detection typically focus on aligning image and text pairs, with an emphasis on out-of-context detection [21, 22, 9, 23, 24, 25]. NewsCLIPpings [18] adopts CLIP-based matching for image-caption congruence. COSMOS [19] employs self-supervised contrastive learning on out-of-context pairs but lacks supervision for fine-grained or localized manipulations. SNIFFER [7] and EMU [17] improve explainability via MLLMs and intent reasoning, yet rely on clean, monolingual captions. HAMMER [9] localizes manipulations through grounding or matching, but does not address subtitle-level or multilingual inconsistencies. MMD-Agent [8] broadens evaluation with mixed-source samples, though without explicit modeling of bilingual or region-specific manipulations. CroMe [26] introduces cross-modal metric learning over image, text, and fused representations, but remains limited to English captions. In contrast, we are the first to detect misinformation involving tampered images and bilingual subtitles, with support for localization, cross-lingual reasoning, and explanation.
@@ 2.2 Multi-modal Large Language Models
Model. Multimodal large language models (MLLMs) [27, 4, 28, 29, 30] integrate vision encoders with pretrained language models to enable unified visual-text reasoning. Flamingo [31] introduces prompt-based multimodal learning, showing strong few-shot results on tasks like VQA. BLIP-2 [28], MiniGPT-4 [29], and LLaVA series [30, 32, 33] further enhance visual instruction tuning, aligning modalities through instruction-following data to support tasks such as VQA, captioning, and dialogue. Recent models like InternVL3 [27], Qwen3-VL [34], and Gemini [35] improve generalization, multilinguality, and open-ended tool use. Despite these advances, most MLLMs focus on grounded understanding or generation and lack supervision for detecting fine-grained cross-modal inconsistencies—critical for misinformation detection [36, 37]. To improve efficiency, many lightweight MLLMs have emerged, such as Gemma [38]; in our work, we adopt Gemma 3 [39] as the base model to balance performance and scalability.
GRPO. Reinforcement learning (RL) has been shown to enhance reasoning in LLMs [40], with DeepSeek-R1 [5, 12] introducing Group Relative Policy Optimization (GRPO) to further improve explanation-driven multimodal performance. GRPO optimizes the relative ranking of outputs within grouped samples, enabling fine-grained discrimination between aligned and misaligned multimodal inputs. However, its potential remains underexplored in misinformation detection tasks involving subtle visual edits and bilingual subtitle inconsistencies.
Prior work [7, 8, 9] rarely tackles misinformation that involves both manipulated images and bilingual subtitles, and often lacks the ability to localize edits, reason across modalities, and adapt to new events. To address these gaps, we propose a unified framework that integrates bilingual visual-text reasoning, manipulation localization, explanation generation, and retrieval-based adaptation.
@@ 3 BiMiBench: A Benchmark for Bilingual Multimodal Misinformation
Existing benchmarks for multimodal misinformation primarily focus on monolingual captions or synthetic mismatches, lacking realism, cross-lingual scope, and fine-grained supervision. In real-world scenarios, misinformation often involves tampered images with bilingual (Chinese-English) subtitles, where inconsistencies may occur in any modality. To address this gap, we introduce BiMiBench, the first benchmark that supports not only bilingual multimodal reasoning but also manipulation, localization, and natural language explanation, enabling comprehensive evaluation in realistic settings.
3.1 Dataset Construction
Details. BiMiBench contains 104,000 samples derived from real-world news content. Specifically, we construct our dataset based on the VisualNews [41] dataset and select the portion sourced from the Washington Post to ensure factual grounding and linguistic consistency. Each sample includes an image paired with Chinese-English subtitles, with misinformation introduced through localized image editing, Chinese subtitle modification, or English subtitle modification. Approximately 80% of the samples contain manipulated content in at least one modality, while the remaining 20% are clean control examples. Image resolutions range from 640×480 to 1024×768 pixels, and subtitles are provided in both English and Chinese to support bilingual misinformation detection. BiMiBench captures cross-modal inconsistencies, offering a challenging testbed for evaluating multimodal and bilingual reasoning capabilities.
@@ 3.1 Dataset Construction
To generate realistic misinformation samples, we design a multi-step data generation pipeline. Let $\mathcal{D}_{\text{orig}}=\{(I_{i},C_{i})\}_{i=1}^{N}$ denote the original dataset from VisualNews, where $I_{i}$ is an image and $C_{i}$ is its associated English caption.
Step 1: Textual Manipulation. Given an image-caption pair $(I_{i},C_{i})$ from VisualNews [41], we use the Gemma 3 [39] model $f_{\text{Gemma}}$ to generate a manipulated caption $\tilde{C}_{i}=f_{\text{Gemma}}(I_{i},C_{i},\text{PROMPT})$, introducing factual inconsistencies while preserving contextual plausibility.
Step 2. Bilingual Translation. We translate the original and manipulated English captions into Chinese using a translation API [42] $\mathcal{T}$, yielding bilingual subtitle pairs: $C^{\text{zh}}_{i}=\mathcal{T}(C_{i}),\tilde{C}^{\text{zh}}_{i}=\mathcal{T}(\tilde{C}_{i})$.
Step 3. Semantic Segmentation. To enable targeted visual editing, we use an instruction-based segmentation model LISA [43]:$f_{\text{LISA}}$ to extract object masks based on the manipulated caption: $M_{i}=f_{\text{LISA}}(I_{i},\tilde{C}_{i},\text{PROMPT})$, where $M_{i}$ highlights regions corresponding to entities in $\tilde{C}_{i}$.
Step 4. Visual Manipulation. We apply a visual manipulation function $\mathcal{V}$ to perform localized editing on the original image using the object masks: $\tilde{I}_{i}=\mathcal{V}(I_{i},M_{i})$, where the edited image $\tilde{I}_{i}$ semantically aligns with the manipulated caption $\tilde{C}_{i}$. To ensure visual diversity and realism, we adopt a mix of recent state-of-the-art image editing techniques, including FLUX [44], VAR [45], and SDXL [46].
@@
All Consistent $I_{i}^{*}=I_{i},C_{i}^{\text{en}}=C_{i},L_{i}^{\text{zh}}=C_{i}^{\text{zh}}$
Image Manipulated $I_{i}^{*}=\tilde{I}_{i},C_{i}^{\text{en}}\neq\tilde{C}_{i}$ or $L_{i}^{\text{zh}}\neq\tilde{C}_{i}^{\text{zh}}$
Both Misaligned $I_{i}^{*}=I_{i},C_{i}^{\text{en}}=\tilde{C}_{i},L_{i}^{\text{zh}}=\tilde{C}_{i}^{\text{zh}}$
Chinese Misaligned $I_{i}^{*}=I_{i},C_{i}^{\text{en}}=C_{i},L_{i}^{\text{zh}}=\tilde{C}_{i}^{\text{zh}}$
English Misaligned $I_{i}^{*}=I_{i},C_{i}^{\text{en}}=\tilde{C}_{i},L_{i}^{\text{zh}}=C_{i}^{\text{zh}}$
All Inconsistent $I_{i}^{*}=\tilde{I}_{i},C_{i}^{\text{en}}=\tilde{C}_{i},L_{i}^{\text{zh}}=\tilde{C}_{i}^{\text{zh}}$
Each sample in BiMiBench is assigned to one of six mutually exclusive categories based on image manipulation and semantic alignment between the image $I_{i}^{*}$, English subtitle $C_{i}^{\text{en}}$, and Chinese subtitle $L_{i}^{\text{zh}}$. The categories cover typical multimodal misinformation patterns: (1) All Consistent: Image and both subtitles are consistent. (2) Image Manipulated: The image is tampered, and at least one subtitle does not align with it. (3) Both Misaligned: Image is real, but both subtitles are misleading. (4) Chinese Misaligned: Image is real, English subtitle is faithful, Chinese is misleading. (5) English Misaligned: Image is real, Chinese subtitle is faithful, English is misleading. (6) All Inconsistent: Image is tampered; both subtitles are misleading. Table 2 provides formal definitions for each class. All samples are manually reviewed for annotation quality and consistency. This design enables structured supervision of complex multimodal misinformation scenarios. Additional examples and detailed analysis of BiMiBench are provided in the Appendix.

### v3

@@ Abstract
The increasing realism of multimodal content has made misinformation more subtle and harder to detect, especially in news media where images are frequently paired with bilingual (e.g., Chinese-English) subtitles. Such content often includes localized image edits and cross-lingual inconsistencies that jointly distort meaning while remaining superficially plausible. We introduce BiMi, a bilingual multimodal framework that jointly performs region-level localization, cross-modal and cross-lingual consistency detection, and natural language explanation for misinformation analysis. To support generalization, BiMi integrates an online retrieval module that supplements model reasoning with up-to-date external context. We further release BiMiBench, a large-scale and comprehensive benchmark constructed by systematically editing real news images and subtitles, comprising 104,000 samples with realistic manipulations across visual and linguistic modalities. To enhance interpretability and cross-modal reasoning consistency, we apply Group Relative Policy Optimization (GRPO) to improve explanation quality, marking the first use of GRPO in this domain. Extensive experiments demonstrate that BiMi outperforms strong baselines by up to +8.9 in classification accuracy, +15.9 in localization accuracy, and +2.5 in explanation BERTScore, advancing state-of-the-art performance in realistic, bilingual misinformation detection. Code, models, and datasets will be released.
@@ 1 Introduction
Figure 1: Comparison of traditional vs. bilingual misinformation detection tasks. Traditional tasks focus on visual-text consistency with limited outputs (left). Our setting uses tampered images and bilingual subtitles, enabling richer outputs including region localization, cross-modal consistency, and explanation (right). Red indicates error, green indicates correctness. Best viewed in color.
The rapid progress of large generative models [achiam2023gpt, rombach2022high, zhao2025surveylargelanguagemodels, qwen, guo2025deepseek, hurst2024gpt] has substantially lowered the barrier to producing highly realistic multimodal content [openai2023dalle3], enabling new creative applications but also heightening the risk of multimodal misinformation, where manipulated images and accompanying text jointly mislead audiences [qi2024sniffer, liu2024mmfakebench]. A recent UNESCO report warns that the widespread adoption of generative-AI tools can be exploited to fabricate convincing visual–textual narratives, thereby creating fertile ground for large-scale multimodal misinformation. Lost in Translation [quelle2025lost, guo2019future] further shows that misinformation frequently crosses language boundaries, underscoring the global risk of cross-lingual diffusion. A striking example occurred in early 2020, when a CDC report confirming the first U.S. case of COVID-19 community transmission was mistranslated on Chinese social media to claim that the virus originated in the United States, fueling public misunderstanding and geopolitical tension. These observations reveal how multimodal misinformation can exploit localized image edits and asymmetric subtitle translations to manipulate public perception across language communities [abdali2024multi]. This creates an urgent need for methods capable of detecting fine-grained multimodal and cross-lingual inconsistencies—capabilities that remain largely underexplored. We therefore formulate the problem as bilingual multimodal misinformation detection: jointly localizing manipulated image regions and identifying cross-lingual inconsistencies while generating faithful natural-language explanations.
Despite notable progress in multimodal misinformation detection [wu2019misinformation, tahmasebi2024multimodal, wan2024dell] and deepfake detection [huang2025so, huang2025sida, zhao2025deepfakebench, shao2025deepfake], existing approaches remain constrained by a predominant focus on coarse-grained image–text alignment and monolingual settings. First, many systems match images and captions only at the global level [qi2024sniffer, liu2024mmfakebench], leaving them unable to localize fine-grained, region-level manipulations such as localized edits or subtle subtitle tampering. Second, their explanation modules typically produce high-level, generic rationales [shao2025deepfake], providing little concrete evidence of why specific content is misleading. Third, no prior work has specifically addressed bilingual subtitle inconsistencies, leaving the semantic divergence introduced by deliberately misleading cross-lingual translations largely unexplored. These limitations collectively underscore the need for a unified framework that can jointly reason over visual content and bilingual text with precise localization and faithful explanation—capabilities that motivate the approach we develop in this work.
@@
However, building such a framework is technically demanding. Achieving fine-grained grounding between image regions and textual cues remains beyond the capability of most existing multimodal large language models (MLLMs). Moreover, producing faithful explanations in a bilingual multimodal setting requires accurate detection and articulation of inconsistencies across both modalities and languages. Cross-lingual reasoning is further complicated by subtle semantic shifts that even fluent translations can introduce, often obscuring misinformation signals. These challenges are especially acute in the news domain, where content evolves rapidly and exhibits high diversity and context dependence, making it non-trivial to design a system that jointly delivers precise localization, robust cross-lingual reasoning, and faithful explanation.
@@
To address these challenges, we introduce BiMi, the first framework that jointly targets bilingual subtitles inconsistency, region-level manipulation localization, and natural-language explanation generation. Figure 1 compares our unified task setup with prior multimodal approaches. To improve generalization to emerging events, BiMi incorporates an online retrieval module that augments model reasoning with real-time external knowledge. We further construct BiMiBench, the first large-scale benchmark for this setting, containing 104,000 news-image samples with realistic manipulations of visual content and bilingual subtitles. Extensive experiments demonstrate that BiMi establishes a new state of the art, outperforming strong baselines by +8.9 in classification accuracy, +15.9 in localization accuracy, and +2.5 in explanation BERTScore, demonstrating robust overall improvements.
Our main contributions are as follows:
@@
  • BiMiBench: the first large-scale benchmark for bilingual multimodal misinformation detection, comprising 104K news–image samples with fine-grained visual manipulations and bilingual subtitle inconsistencies.
@@
  • BiMi Framework: a unified model that detects image–subtitle misinformation through region-level manipulation localization, multimodal and cross-lingual consistency detection, and natural-language explanation generation, directly addressing key limitations of prior work.
@@
  • GRPO for Explanation: the first application of Group Relative Policy Optimization (GRPO) [shao2024deepseekmath] to improve the quality and faithfulness of bilingual multimodal explanations.
@@
  • State-of-the-Art Results: BiMi achieves state-of-the-art performance on BiMiBench, with significant improvements in classification, localization, and explanation BERTScore over strong baselines.
@@ 2.1 Misinformation Detection
Datasets. Early misinformation detection datasets can be grouped into three main categories. Text-only: FEVER [Thorne18Fever] and FakeNewsNet [shu2020fakenewsnet] target textual fact verification and lack visual modality. Basic multimodal: Fakeddit [nakamura2019r], MAIM [jaiswal2017multimedia], and EMU [da-etal-2021-edited] pair images with text but are monolingual, use coarse edits, and provide no supervision for localization or explanation. Fine-grained multimodal: NewsCLIPpings [luo-etal-2021-newsclippings], COSMOS [aneja2021cosmos], MMFakeBench [liu2024mmfakebench], and DGM4 [shao2023detecting] enable richer reasoning yet still focus on caption-style inputs and ignore subtitle-level or cross-lingual inconsistencies. We introduce BiMiBench, which covers five types of visual and textual misinformation, uniquely supporting bilingual subtitles, region-level localization, and dual-modality tampering (Table 1); a perceptual analysis (Figure 2) shows it achieves higher realism and diversity than existing datasets.
Table 1: Comparison of misinformation datasets. BiMiBench uniquely supports bilingual and localized manipulations.
Dataset Sub. Bi. Text Vis. Loc. News Mask
FEVER [Thorne18Fever]
FakeNewsNet [shu2020fakenewsnet]
Fakeddit [nakamura2019r]
MAIM [jaiswal2017multimedia]
EMU [da-etal-2021-edited]
NewsCLIPpings [luo-etal-2021-newsclippings]
COSMOS [aneja2021cosmos]
DGM4 [shao2023detecting]
MMFakeBench [liu2024mmfakebench]
@@
Methods. Current multimodal misinformation methods mainly align image–text pairs for out-of-context detection [przybyla2020capturing, qi2021improving, shao2023detecting]. NewsCLIPpings [luo-etal-2021-newsclippings] and COSMOS [aneja2021cosmos] use CLIP-based or contrastive learning but lack supervision for fine-grained localization. SNIFFER [qi2024sniffer] and EMU [da-etal-2021-edited] add explainability via MLLMs yet rely on clean monolingual captions. HAMMER [shao2023detecting] localizes manipulations but ignores subtitle-level or multilingual cues. MMD-Agent [liu2024mmfakebench] and CroMe [choi2025cromemultimodalfakenews] broaden evaluation but remain limited to English captions. In contrast, our framework is the first to handle tampered images with bilingual subtitles, supporting localization, reasoning and explanation.
@@ 2.2 Multi-modal Large Language Models
Model. Multimodal large language models (MLLMs) have greatly advanced cross-modal reasoning. Representative models such as Flamingo [alayrac2022flamingo], BLIP-2 [li2023blip], MiniGPT-4 [zhu2023minigpt4enhancingvisionlanguageunderstanding], Gemini [gemini2023, comanici2025gemini], InternVL-3 [zhu2025internvl3exploringadvancedtraining], Qwen-VL [bai2023qwenvlversatilevisionlanguagemodel, xu2025qwen2, qwen3], Gpt-4o [hurst2024gpt] and LLaVA [liu2023llava, liu2024llavanext] extend pretrained LLMs with cross-modal attention and instruction tuning for vision-language tasks. DeepSeek-R1 [guo2025deepseek, shao2024deepseekmath] further employs GRPO to enhance explanation quality. Yet most MLLMs focus on grounded understanding or generation [wu2025grounded, lin2025boosting] and are not trained to detect cross-modal inconsistencies. We adapt Gemma 3 [team2025gemma] to strengthen reasoning on visual and bilingual cues for misinformation detection.
Training. MLLMs training typically includes large-scale pre-training followed by post-training with supervised fine-tuning (SFT) and reinforcement learning from human feedback (RLHF)[ouyang2022training]. RLHF aligns outputs with human preferences using a reward model; methods such as Proximal Policy Optimization (PPO) [schulman2017proximal], Direct Preference Optimization (DPO) [rafailov2023direct], and GRPO [guo2025deepseek, shao2024deepseekmath] refine policies through preference ranking. RLHF remains explored for detecting visual edits and bilingual subtitle inconsistencies.
Figure 2: Image quality comparison across datasets using the perceptual evaluation method of [chen2024evaluating]. BiMiBench achieves higher visual quality across dimensions than prior datasets, indicating closer alignment with real-world social media imagery.
@@ 3 BiMiBench: A Benchmark for Bilingual Multimodal Misinformation
Existing benchmarks for multimodal misinformation primarily focus on monolingual captions or synthetic mismatches, lacking realism, cross-lingual scope, and fine-grained supervision. In real-world scenarios, misinformation often involves tampered images with bilingual (Chinese-English) subtitles, where inconsistencies may occur in any modality. We introduce BiMiBench, a benchmark for multimodal misinformation detection with bilingual subtitles, providing labels and explanations for joint evaluation of classification, localization, and natural language explanation quality.
3.1 Benchmark Construction
Details. BiMiBench comprises 104,000 real-world news samples derived from the VisualNews [liu-etal-2021-visual] corpus, a professionally curated collection of image–text pairs with broad topical diversity and reliable editorial quality. Each sample pairs an image with Chinese–English subtitles; misinformation is introduced via localized image edits or Chinese/English subtitle modifications. About 80% of the samples contain manipulations in at least one modality and 20% serve as clean controls. Images (640×480–1024×768) and bilingual subtitles provide a challenging testbed for evaluating multimodal and bilingual reasoning.
@@ 3.1 Benchmark Construction
To generate realistic misinformation samples, we design a multi-step data generation pipeline shown in Figure 3. Let $\mathcal{D}_{\text{orig}}=\{(I_{i},C_{i})\}_{i=1}^{N}$ denote the original dataset, where $I_{i}$ is an image and $C_{i}$ is its associated English caption.
Step 1: Textual Manipulation. Given an image-caption pair $(I_{i},C_{i})$ from VisualNews [liu-etal-2021-visual], we use the Gemma 3 [team2025gemma] model $f_{\text{Gemma}}$ to generate a manipulated caption $\tilde{C}_{i}=f_{\text{Gemma}}(I_{i},C_{i},\text{PROMPT})$, introducing inconsistencies while preserving contextual plausibility.
Step 2. Bilingual Translation. We translate the original and manipulated English captions into Chinese using a translation API [googleapi2024] $\mathcal{T}$, selected for its accuracy and reliability in large-scale bilingual news translation, yielding bilingual subtitle pairs: $C^{\text{zh}}_{i}=\mathcal{T}(C_{i}),\tilde{C}^{\text{zh}}_{i}=\mathcal{T}(\tilde{C}_{i})$.
[…]
Step 4. Visual Manipulation. We apply a visual manipulation function $\mathcal{V}$ to perform localized editing on the original image using the object masks: $\tilde{I}_{i}=\mathcal{V}(I_{i},M_{i})$, where the edited image $\tilde{I}_{i}$ semantically aligns with the manipulated caption $\tilde{C}_{i}$. To ensure visual diversity and realism, we adopt a mix of recent state-of-the-art image editing techniques, including FLUX [flux2024], VAR [tian2024visual], and SDXL [podell2023sdxlimprovinglatentdiffusion].
@@
All Consistent $I_{i}^{*}=I_{i},\,C_{i}^{\mathrm{en}}=C_{i},\,L_{i}^{\mathrm{zh}}=C_{i}^{\mathrm{zh}}$
Image Manipulated $I_{i}^{*}=\tilde{I}_{i},\,C_{i}^{\mathrm{en}}\neq\tilde{C}_{i}$ or $L_{i}^{\mathrm{zh}}\neq\tilde{C}_{i}^{\mathrm{zh}}$
Both Misaligned $I_{i}^{*}=I_{i},\,C_{i}^{\mathrm{en}}=\tilde{C}_{i},\,L_{i}^{\mathrm{zh}}=\tilde{C}_{i}^{\mathrm{zh}}$
Chinese Misaligned $I_{i}^{*}=I_{i},\,C_{i}^{\mathrm{en}}=C_{i},\,L_{i}^{\mathrm{zh}}=\tilde{C}_{i}^{\mathrm{zh}}$
English Misaligned $I_{i}^{*}=I_{i},\,C_{i}^{\mathrm{en}}=\tilde{C}_{i},\,L_{i}^{\mathrm{zh}}=C_{i}^{\mathrm{zh}}$
All Inconsistent $I_{i}^{*}=\tilde{I}_{i},\,C_{i}^{\mathrm{en}}=\tilde{C}_{i},\,L_{i}^{\mathrm{zh}}=\tilde{C}_{i}^{\mathrm{zh}}$
Categories. Each BiMiBench sample is labeled into one of six categories according to consistency among image $I_{i}^{*}$, English subtitle $C_{i}^{\text{en}}$, and Chinese subtitle $L_{i}^{\text{zh}}$. (1) All Consistent: all modalities agree; (2) Image Manipulated: image tampered, at least one subtitle mismatched; (3) Both Misaligned: image real, both subtitles misleading; (4) Chinese Misaligned: only Chinese subtitle misleading; (5) English Misaligned: only English subtitle misleading; (6) All Inconsistent: image tampered and both subtitles misleading. Formal definitions show in Table 2.
All samples were manually reviewed to ensure annotation quality and consistency. Each item was independently checked by two annotators with a senior reviewer resolving disagreements, following predefined guidelines on factual correctness and translation fidelity. To address potential ethical and privacy concerns, we used only publicly available news content, removed any personally identifiable information, and release the dataset solely for non-commercial research in accordance with the source licenses. This design enables structured supervision of complex multimodal misinformation scenarios while maintaining high standards of data integrity and ethical compliance. Detailed information about the BiMiBench can be found in Appendix.